OpenStudy (anonymous):

I don't understand .. :( The heights in inches for five student are: 60, 67, 64, 54, 58 How would the median change if a student 68 inches tall is added to the group? increased by 2 decreased by 1 increased by 1 no change

OpenStudy (lgbasallote):

first arrange the given in increasing order 54, 58, 60, 64, 67 right?

OpenStudy (anonymous):

Yes ...?

OpenStudy (lgbasallote):

so what is the median there?

OpenStudy (anonymous):

60 ..

OpenStudy (lgbasallote):

right...now if you add 68 to the given you'll have 54, 58, 60, 64, 67, 68 what is the median this time?

OpenStudy (anonymous):

.... 62..?

OpenStudy (lgbasallote):

right..

OpenStudy (lgbasallote):

so what's the difference from 60 to 62

OpenStudy (anonymous):

it increases by two

OpenStudy (lgbasallote):

right!

OpenStudy (anonymous):

Yaaay :D Its so much easier when some one explains it step by step .. Thankyou(:

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

Will you help me with . one .. more problem ... ? Pleeeeeeeeease .. :D

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

i Do not understand this one either ... What is the approximate distance from (4, 6) to (-2,1) ?

OpenStudy (lgbasallote):

you're going to use the distance formula on that one \[\huge \text{distance} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

i dont know how to use it ... would it be .. ( 4 - -2 ) ^2 + ( 6 - 1 ) ^2 ..??

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

note: 4 - -2 will become 4 + 2

OpenStudy (anonymous):

Okayy .. Hold on let me do the math .. lol

OpenStudy (anonymous):

4+2 = 6 ^2 = 36 ..? + 6 -1 = 5^2 = 25 ? .. so ... 36 + 25 ? = 61 .. and the square root is ..? .. i dont understand :(( ughhhhhh

OpenStudy (lgbasallote):

you're doing fine

OpenStudy (anonymous):

7.8 ..

OpenStudy (anonymous):

and the choices are ... 5 units 9 units 6 units 8 units ... so it would be 8 units..?

OpenStudy (anonymous):

...? yes

OpenStudy (lgbasallote):

i think so..yeah

OpenStudy (anonymous):

Oh okay .. lol :D